OpenStudy (anonymous):

On a meter stick acting like a seesaw. The mass m1 is 0.53 kg and it is located at x1 = 25 cm. The pivot point is represented by the solid triangle located at x = 60 cm. The mass of the meter stick (mms = 0.40 kg) is located at its geometric center, xms = 50 cm. The mass m2 is 0.16 kg and it is located at x2 = 85 cm. Calculate the net torque (in N⋅m with the proper sign) due to these three weights. Use g = 9.79 m/s2.

OpenStudy (anonymous):

assuming the meterstick is horizontal, let cw torque be positive, then\[\tau _{net} = -m _{1}g \times (60cm - 25cm) - m _{ms} g \times (60cm - 50cm) + m _{2}g \times (85cm - 60cm)\]plug in the masses and g

OpenStudy (anonymous):

...and convert cm to m